Consider the following supply and demand curves for a certain product.

QS = 25,000P

QD = 50,000 - 10,000P

a. Plot the demand and supply curves.

b. What are the equilibrium price and equilibrium quantity for the industry? Determine

the answer both algebraically and graphically. (Round to the nearest cent.)



1
Expert's answer
2021-02-11T17:12:21-0500

a) Let's plot the demand and supply curves in Excel:



b) Let's solve algebraically:


"Q_s=Q_d,""25000P=50000-10000P,""P=\$1.43"

Then,

"Q=50000-10000\cdot\$1.43=35700."

As we can see from the graph, "P=\$1.43" and "Q=35700".